A data set has nine values. The mean of the set is five. When a tenth value is added, the mean becomes six. What is the tenth value?
What is the sum of the original nine data value
Make a Plan Write and solve an equation to find the tenth value
But is you can help me its in math but that if your a math major but its okay if you aren't i have tryed to do this problem more than once.

Answer by stanbon(57967)

A data set has nine values. The mean of the set is five.
Total points was 9*5 = 45
--------------------------------------
When a tenth value is added, the mean becomes six. What is the tenth value?
Total points will be 10*6 = 60 points.
--
The 10th value must have been 60-45 = 15
------------------------------------------------
What is the sum of the original nine data values : 45
----
Cheers,
Stan H.
==========